OpenStudy (anonymous):

Find the equilibria of the system dx/dt = 10x(1-x/10)-20xy and dy/dt = -5y+xy/20 @amistre64 @jim_thompson5910 @Mertsj @ganeshie8 @Callisto @zepdrix @radar @Luigi0210 @whpalmer4 @Compassionate @wio @e.mccormick @shamil98 @mukushla @timo86m @nincompoop @hba @beccaboo333 @Zale101 @Preetha @bookworm00981 @wolfe8

OpenStudy (nincompoop):

I found out that I can actually do DE's

OpenStudy (anonymous):

I got equilibria (0,0) , (10/0) and (1000, -99/2) but I think they are wrong for some reason

OpenStudy (beccaboo333):

I cannot help you. Please stop tagging me and so many people at once.

OpenStudy (nincompoop):

inky can help you

OpenStudy (inkyvoyd):

inky is a hedgehog :D

OpenStudy (nincompoop):

y'= -5y+xy/20 ? is this all over 20 or just the xy part?

OpenStudy (anonymous):

Just the xy part

OpenStudy (agent47):

I'm confused.. but what I would do is.. dy/dx=dy/dt/dx/dt and integrate

OpenStudy (tkhunny):

Dear FMP, Haven't you already been temporarily banned for gross over-tagging? I have a plan. Quit it. It just is not necessary and is quite annoying. Seriously, stop it.

OpenStudy (anonymous):

Well answer my question then, noob.

OpenStudy (tkhunny):

Definitely seeing (0,0) and (10,0). Where did you get the other one? There is some argument around x = 100. But are you learning anything? This is the question. If you just come back and do the same thing over again, it seems like you are not learning. I recommend learning. Give it a go.

OpenStudy (anonymous):

Well I got a 108/100 on the first exam so I must be.

OpenStudy (anonymous):

From the first equation you get 1-x/10-2y=0 so you get x = 10/20y then you sub it to the 2nd equation to get y=-99/2 then you sub this back into the irinal equation to get x=1000

OpenStudy (tkhunny):

I wasn't talking about your mathematics. I happen to have infinite patience. Not everyone does. Your general demeanor could use some learning. Your "1" seems magically to have disappeared. Please try that "first equation" again.

OpenStudy (anonymous):

I'm not so sure about that 1000. The other two are right though.

OpenStudy (tkhunny):

The x = 1000 doesn't exist. It is just an algebra error.